# Physics-Informed Inductive Biases for Voltage Prediction in Distribution Grids

### v1

@@ 4.1 Baseline Model
When predicting voltage magnitude on known grid configurations, the baseline GNN attains an average RMSE of $\mathbf{0.0067}\,\mathrm{p.u.}$ for voltage magnitude, and $\mathbf{0.1573^{\circ}}$ for voltage angle. For the OOD experiments, the baseline yields a mean performance of $\overline{\mathrm{RMSE}}(|V|)=\mathbf{0.0179}\,\mathrm{p.u.}$ and $\overline{\mathrm{RMSE}}(\angle V)=\mathbf{1.3271^{\circ}}$. Notably, this model proves to be quite stable in its OOD prediction of voltage magnitude. Mean training time is 1187s over all experiments.
@@ 4.2 Complex-Valued Learning
The complex-valued model excels on both magnitude and angle prediction in across test cases. It achieves $\mathbf{0.00239}^{\circ}$ angle RMSE and $\mathbf{0.00555}$ p.u. magnitude RMSE when predicting on known grid configurations, proving to be the best among all methods. Across the generalization experiments, we observe a continued stable accuracy, achieving $\overline{\mathrm{RMSE}}(\angle V)=\mathbf{0.0212}^{\circ}$ and $\overline{\mathrm{RMSE}}(|V|)=\mathbf{0.0189}$ p.u. These results underscore the remarkable angle predictive performance of this model for both familiar and unseen networks. However, while the angle predictive accuracy is the best among the variants, the complex-valued model does not achieve the top result for voltage magnitude prediction. Additionally, this performance gain comes with the highest computational burden, producing an average training time of 4346s across all experiments.
@@ 4.3 Physics-Informed Learning
The supervised physics-informed variant exhibits an increase in performance compared to the baseline model for almost all experiments, illustrating the impact of this PI fine-tuning mechanism. For the known grids setting, it reaches $\text{RMSE}(|V|)=\mathbf{0.00638}$ p.u. and $\text{RMSE}(\angle V)=\mathbf{0.1488^{\circ}}$. Across OOD benchmarks, the method exhibits the most accurate and stable performance for voltage magnitude prediction, with $\overline{\mathrm{RMSE}}(|V|)=\mathbf{0.0166}$ p.u., and the smallest standard deviation. When considering voltage angle, its mean performance ($\overline{\mathrm{RMSE}}(\angle V)=\mathbf{1.263^{\circ}}$) is only outperformed by the complex-valued model. The mean training time is 1283s across all experiments.
@@ 4.4 Residual Learning
The residual approach provides performance comparable to the baseline model. When predicting on familiar grids, it attains $\text{RMSE}(|V|)=\mathbf{0.00662}$ p.u. and $\text{RMSE}(\angle V)=\mathbf{0.1573^{\circ}}$, showing nearly identical performance to the baseline. Across generalization experiments, accuracy and stability are also similar to the baseline, with $\overline{\mathrm{RMSE}}(|V|)=\mathbf{0.01828}$ p.u. and $\overline{\mathrm{RMSE}}(\angle V)=\mathbf{1.342^{\circ}}$. However, it is worth noting that, unlike the baseline, the residuals model proved to be the best performing model in certain situations, as made evident in the LV-semiurb5 and LV-urban6 OOD experiments. Training times are comparable to the baseline (1200s).
@@ 5.1 Inductive Biases for Voltage Prediction
When estimating voltage magnitude and angle for networks similar to those seen in training, the complex-valued model provides the most accurate predictions. Voltage angle is particularly impressive, significantly outperforming the other variants by approximately two orders of magnitude. Generally, all physics-informed inductive biases seem to improve the results of the already-performant baseline model when they have had prior exposure to all grid types. This demonstrates that all three inductive biases can be used to assist in data-driven learning of power grid characteristics.
When looking at the results of the generalization experiments, the presence of unseen grid topologies clearly influences model accuracy, with average performance dropping by one order of magnitude on average. In this setting, the physics-informed variant attains the best performance on average for voltage magnitude and the lowest variability across individual OOD grids. The complex-valued model continues to thrive in voltage angle prediction. In fact, the complex-valued model’s results on the OOD tests greatly surpass those of the other models under known grid configurations, highlighting its fidelity for voltage angle prediction.
@@ 5.2 Real-world application
As an additional interpretation, we compare the best performing models for voltage prediction to a commonly used solving method for voltage prediction: DC Power Flow. The DC power flow approximation linearizes the full AC method such that a non-iterative, convergent solution can be found. However, this simplification assumes small angle differences and voltages close to nominal, and thus struggles in situations with low voltage drops. Looking at Figure 2, we see that using GNN-based solvers, we are able to match the performance of DC PF for both known and unseen grids by adding physics informed loss functions for model fine-tuning. Additionally, by using complex-valued neural networks, we are able to outperform DC PF by one order of magnitude in voltage angle prediction. This is a significant result because it improves on a commonly observed limitation that GNN-based power flow methods perform well on known-configurations but are ineffective in predicting PF from unseen grids during training, and are usually outperformed by simplified models such as DC PF (Hansen et al., 2023; Yaniv et al., 2023; Okoyomon & Goebel, 2025).
@@
Figure 2: Comparison of Best-Performing models to DC PF.
@@ 5.3 Conclusion and Future Work
This study evaluates three inductive-bias strategies for voltage prediction in distribution grids. Using a dataset of heterogeneous low- and medium-voltage networks, we assess performance across familiar topologies and out-of-distribution evaluation cases. Our findings highlight complementary strengths. The complex-valued model achieves state-of-the-art angle prediction for both known and unknown grids, surpassing even DC power flow approximations, though at higher computational cost. The physics-informed variant consistently improves magnitude prediction and exhibits the lowest variability across unseen topologies, demonstrating robustness rooted in physical consistency. In contrast, the residual learning approach provided limited overall gains, though it excelled in select cases. Collectively, these results show that carefully chosen inductive biases substantially enhance the reliability of machine learning models for power system analysis.
Future research should extend these insights along three directions. Firstly, a deeper examination of physics-constrained training regimes is needed, including theoretical analysis and open benchmarking, to clarify when physics-informed losses improve generalization. For our particular study, this variant was the most sensitive to hyperparameters and required significant knowledge of the pandapower modeling library to produce an effective solution. Even then, the physics-informed loss could not be used competitively in a purely unsupervised manor (nor as a weighted loss addition), with performance results two orders of magnitude worse than the other variants at best. This finding comes at no surprise when considering the non-linear, non-convex nature of the AC power balance equations, but is nonetheless in conflict with several other works that tout its success. Secondly, evaluation on non-European grid datasets (e.g. Smart-DS) would test the applicability of these methods in unbalanced, three-phase networks with different structural characteristics. Finally, exploring hybrid models that combine multiple inductive biases could yield situation-dependent predictors, though early results suggest such benefits are non-additive and grid-specific.
@@ Appendix A Complete Benchmarking Results
We present the full benchmarking results in Table 2, which provides the raw performance metrics for all models across all testing grids, using RMSE, MAPE and training time as evaluation criteria.

### v2

@@ 4.1 Baseline Model
When predicting voltage magnitude on known grid configurations, the baseline GNN attains an average RMSE of $\mathbf{0.0067}\,\mathrm{p.u.}$ for voltage magnitude, and $\mathbf{0.1573^{\circ}}$ for voltage angle. For the OOD experiments, the baseline yields a mean performance of $\overline{\mathrm{RMSE}}(|V|)=\mathbf{0.0179}\,\mathrm{p.u.}$ and $\overline{\mathrm{RMSE}}(\angle V)=\mathbf{1.3271^{\circ}}$. Notably, this model proves to be quite stable in its OOD prediction of voltage magnitude. The model contains 302.7K parameters and its training and inference times are approximately 1200s and 2ms for all experiments.
@@ 4.2 Complex-Valued Learning
The complex-valued model excels on both magnitude and angle prediction across test cases. It achieves $\mathbf{0.00239}^{\circ}$ angle RMSE and $\mathbf{0.00555}$ p.u. magnitude RMSE when predicting on known grid configurations, proving to be the best among all methods. Across the generalization experiments, we observe a very tight distribution for both voltage magnitude and angle. One bad performance outlier drives the $\overline{\mathrm{RMSE}}(|V|)$ up to $\mathbf{0.0715}$ p.u, while $\overline{\mathrm{RMSE}}(\angle V)$ remains stable at $\mathbf{0.0212}^{\circ}$. These results underscore the remarkable angle predictive performance of this model for both familiar and unseen networks. This performance gain comes with the highest computational burden, producing an average of approximately 4x training time, 3x inference time, and 2x model capacity compared to the other variants.
@@ 4.3 Physics-Informed Learning
The supervised physics-informed variant exhibits an increase in performance compared to the baseline model for almost all experiments, illustrating the impact of this PI fine-tuning mechanism. For the known grids setting, it reaches $\text{RMSE}(|V|)=\mathbf{0.00638}$ p.u. and $\text{RMSE}(\angle V)=\mathbf{0.1488^{\circ}}$. Across OOD benchmarks, the method exhibits the most accurate and stable performance for voltage magnitude prediction, with $\overline{\mathrm{RMSE}}(|V|)=\mathbf{0.0166}$ p.u., and the smallest standard deviation. When considering voltage angle, its mean performance ($\overline{\mathrm{RMSE}}(\angle V)=\mathbf{1.263^{\circ}}$) is only outperformed by the complex-valued model. Model efficiency metrics are very similar to the baseline model, with a slight increase in training time ($\approx 100s$) due to the physics-informed fine-tuning step.
@@ 4.4 Residual Learning
The residual approach provides performance comparable to the baseline model. When predicting on familiar grids, it attains $\text{RMSE}(|V|)=\mathbf{0.00662}$ p.u. and $\text{RMSE}(\angle V)=\mathbf{0.1573^{\circ}}$, showing nearly identical performance to the baseline. Across generalization experiments, accuracy and stability are also similar to the baseline, with $\overline{\mathrm{RMSE}}(|V|)=\mathbf{0.01828}$ p.u. and $\overline{\mathrm{RMSE}}(\angle V)=\mathbf{1.342^{\circ}}$. However, it is worth noting that, unlike the baseline, the residuals model proved to be the best performing model in certain situations, as made evident in the LV-semiurb5 and LV-urban6 OOD experiments. Model efficiencies are comparable to the baseline.
(a) Voltage Prediction
(b) Model Efficiency
Figure 2: Summary of physics-informed model benchmarking metrics across all experiments
@@ 5.1 Inductive Biases for Voltage Prediction
When estimating voltage magnitude and angle for networks similar to those seen in training, our new complex-valued model provides the most accurate predictions. Voltage angle is particularly impressive, significantly outperforming the other variants by approximately two orders of magnitude. Generally, all physics-informed inductive biases seem to improve the results of the already-performant baseline model for the known grids experiment case. This demonstrates that all three inductive biases can be used to assist in data-driven learning of power grid characteristics.
When looking at the results of the generalization experiments, the presence of unseen grid topologies clearly influences model accuracy, with average performance dropping by one order of magnitude on average. In this setting, the physics-informed variant attains the best performance on average for voltage magnitude and the lowest variability across individual OOD grids, though the improvements are minor. The complex-valued model continues to thrive in voltage angle prediction. In fact, the complex-valued model’s results on the OOD tests greatly surpass even the predictions of other models under known grid configurations, highlighting its fidelity for voltage angle prediction.
@@ 5.2 Real-world application
Over known topologies, all models achieve a voltage magnitude RMSE of roughly $0.5\%$, matching the accuracy of iMSys household meters (EFR GmbH, 2022) and aligning with the practical noise floor of commercially deployed measurement hardware. Voltage angle errors are also well below the $1^{\circ}$ metering standard, reaching RMSE values up to three orders of magnitude below this threshold. These findings support the use of such models for distribution grid operation in familiar networks, as they can safely be deployed with limited uncertainty. However, under the OOD testing, voltage angle errors increase by one order of magnitude and only the complex-valued model reliably stays under this $1^{\circ}$ threshold over all cases. A similar observation can be made for voltage magnitude prediction, as even the best performing model produces errors that are higher than the 0.5% measurement standard by one order of magnitude. This highlights the significance of comprehensive generalization studies for robust assessment of machine learning in distribution grids and represents a key finding for establishing practically viable solutions for real-world deployment. Unlike previous works that primarily assess performance on seen or similarly sized topologies, we systematically show that state of the art GNN architectures do not attain the necessary level of accuracy in OOD scenarios. While physics-informed inductive biases may improve the generalization capability, only for the complex valued model did performance stay below the desired threshold for voltage angle, while all other variants exceed acceptable error margins. This indicates the need for a shift in focus in machine learning methodology from standard predictive accuracy to generalization performance assessment in order to take meaningful steps towards real-world application.
As an additional interpretation, we compare the best performing models for voltage prediction to a commonly used simplification of the full AC Power Flow solution: DC Power Flow. The DC power flow approximation linearizes the full AC method such that a non-iterative, convergent solution can be found. However, this simplification ignores reactive power and assumes small angle differences and a fixed voltage magnitude of 1.0 p.u. for all buses, and thus often leads to inaccurate predictions in situations with larger voltage drops. We present this baseline not due to its sophistication but rather due to its relevance as an empirically observed baseline for generalization failure, as established in prior benchmarking studies (Hansen et al., 2022; Yaniv et al., 2023; Okoyomon & Goebel, 2025). These works commonly observe the limitation that GNN-based power flow methods perform well on known-configurations but are ineffective in predicting PF from unseen grids during training, and are usually outperformed by simplified models such as DC PF. Looking at Figure 3, we see that our PFLoss GNN solver is able to significantly outperform DCPF for known networks and match its performance for unseen grids. Additionally, by using complex-valued neural networks, we are able to outperform DC PF by one order of magnitude in voltage angle prediction across all test cases. These results emphasize the promise of physics-informed models as a robust replacement for existing analytical models. For a secondary analytical baseline, we compare performance results to the well-established LinDistFlow method (Baran & Wu, 1989) in Appendix D.2.
@@
Figure 3: Comparison of Best-Performing models to DC PF
@@ 5.3 Conclusion and Future Work
This study evaluates three inductive-bias strategies for voltage prediction in distribution grids. Using a dataset of heterogeneous low- and medium-voltage networks, we assess performance across familiar topologies and out-of-distribution cases. Our findings highlight complementary strengths. Our novel complex-valued model achieves state-of-the-art angle prediction for both known and unknown grids, surpassing all other approximations, though at higher computational cost. The physics-informed variant consistently improves magnitude prediction and exhibits the lowest variability across unseen topologies, demonstrating robustness rooted in physical consistency, but this is only a slight improvement above the strong baseline. In contrast, the residual learning approach provided limited overall gains, though it excelled in select cases and seems to be able to be synergetic with other methods. Collectively, these results show that carefully chosen inductive biases substantially enhance the reliability of machine learning models for power system analysis.
Future research should extend these insights along three directions. Firstly, a deeper examination of physics-constrained training regimes is needed, including theoretical analysis and open benchmarking, to clarify when physics-informed losses improve generalization. For our particular study, this variant was the most sensitive to hyperparameters and required significant knowledge of the pandapower modeling library to produce an effective solution. Even then, the physics-informed loss could not be used competitively in a purely unsupervised manor, with performance results two orders of magnitude worse than the other variants at best. This finding comes at no surprise when considering the non-linear, non-convex nature of the AC power balance equations, but is nonetheless in conflict with several other works that tout its success. Secondly, evaluation on non-European grid datasets (e.g. Smart-DS) would test the applicability of these methods in unbalanced, three-phase networks with different structural characteristics. Finally, exploring hybrid models that combine multiple inductive biases could yield situation-dependent predictors, though early results suggest that such benefits are non-additive and grid-specific (Appendix B.1).
@@ Appendix A Complete Benchmarking Results
We present the full benchmarking results in Table 2, which provides the raw performance metrics for all inductive bias models across all testing grids, using RMSE, training time, inference time, and model capacity as evaluation criteria.